OpenStudy (anonymous):

What power of an engine is required to pump 2450 N of water per second from a well 50 m deep to the surface? (a) 2.45*10^3 watts (b) 2.45*10^4 watts (c) 2.45*10^5 watts (d) none of these

OpenStudy (anonymous):

@Michele_Laino please can you solve it?

OpenStudy (anonymous):

okay so work is Fd so the answer is 2450 N times 50 m

OpenStudy (anonymous):

wait... per second

OpenStudy (anonymous):

ignore my last answer

OpenStudy (anonymous):

I think you're right because a watt is a N-m/s, so the units work out

OpenStudy (michele_laino):

we have to consider the manometric head H, which is about 50 m, so the requested power P is: \[\Large P = F\left( {H + h} \right) = 2450\left( {50 + 50} \right) = ...?\] being h the deepness of the well

OpenStudy (michele_laino):

sorry H is equal to the ratio between the pressure difference and the specific weight of the water and H+h is the manometric head or manometric prevalence

OpenStudy (anonymous):

would you please tell me a little about manometric head,50 m deepness is okay to understand but what about additional 50m that you added

OpenStudy (michele_laino):

since the formula for computing the power, is: \[\large P = \frac{{FH}}{t}\] where H is the total prevalence, whose formula is: \[H = h + \frac{{\Delta p}}{\gamma }\] \Delta p is the difference between the pressures

OpenStudy (michele_laino):



OpenStudy (michele_laino):